Mrs. West is 14 years younger than her aunt. If Mrs. West's age in years is as much below 60 as her aunt's age is over 40, how old is each?
Answer by ptaylor(2198)

Let x=aunt's age
then x-14=Mrs West's age
so we have:
60-(x-14)=x-40 simplify
60-x+14=x-40 subtract x from each side and also subtract 60 and 14 from each side
60-60-x-x+14-14=x-x-40-60-14 collect like terms
-2x=-114
x=57------------aunt's age
x-14=57-14= 43----------Mrs West's age
ck
60-43=57-40
17=17
Hope this helps---ptaylor